# A note on cyclic non-MDS matrices

### v1

@@ 1 Introduction
Symmetric key primitives, like block ciphers and hash functions, rely on various components that provide “confusion and diffusion”. These concepts were introduced by Claude Shannon in $1949$, in his seminal paper “Communication Theory of Secrecy System” [31]. Confusion and diffusion layers play a pivotal role in the security and efficiency of cryptographic schemes based on block ciphers and hash functions. The goal of the confusion layer is to hide the statistical relationship between the ciphertext and message whereas, the diffusion layer guarantees that every bit of the message and secret key affects multiple bits of the ciphertext, ensuring that, after several rounds, all output bits depend on every input bit. Therefore, the choice of a diffusion layer plays an important role in providing security against differential cryptanalysis [4] and linear cryptanalysis [26] which are the two most important cryptanalysis of block ciphers. The strength of a diffusion layer is usually assessed by its branch number. As a result, constructing diffusion layers with higher branch numbers and low-cost implementations is a challenge for cipher design. Consequently, many block ciphers use Maximum Distance Separable (MDS) matrices, known for their optimal branch numbers, in their diffusion layers. These matrices are redundant parts of systematic generator matrices of MDS codes and used in popular ciphers like AES [12], LED [17], SQUARE [11], SHARK [33] as well as hash functions such as Whirlpool [3] and PHOTON [16].
The construction of diffusion layer using MDS matrices primarily employs two main approaches: recursive and non-recursive. In recursive construction, a companion or sparse matrix $A$ of order $n\times n$ is used such that $A^{n}$ is an MDS matrix. Block cipher LED [17], and hash function PHOTON [16] used this recursive based construction. In contrast, non-recursive methods encompass various techniques, such as the use of Cauchy matrices [18, 25], Vandermonde matrices [23, 30], Hadamard matrices [32], and their generalizations [27]. Furthermore, inspired by the use of circulant MDS matrix in the diffusion layer of AES, various authors [5, 19, 20, 24] have explored the construction of circulant MDS matrices. In [20], the authors extended circulant matrices to circulant-like and generalized circulant MDS matrices.
@@
In search of orthogonal MDS matrices, Liu and Sim introduced cyclic matrices as a generalization of circulant matrices [24]. They proved that $2^{d}\times 2^{d}$ left-circulant MDS matrices, a subclass of cyclic matrices, cannot be orthogonal. Furthermore, in [5], the authors introduced $\theta$-circulant matrices over the quasi polynomial ring. Later, in $2022$, Adhiguna et al. studied the orthogonal property of $\theta$-circulant MDS matrices in [2] over the quasi polynomial ring. Several other authors [7, 14] extended the search for MDS matrices over rings and modules. The authors of [28, 29] studied orthogonal MDS matrices over the Galois ring $GR(2^{n},k)$.
@@ 3 Notations and Preliminaries
Let $\mathbb{F}_{q}$ denote the finite field with $q=p^{m}$ elements where $p$ is a prime number and $m$ is a positive integer. Consider an $[n,k,d]$ linear error-correcting code $\mathcal{C}$ over $\mathbb{F}_{q}$. This code $\mathcal{C}$ is a subspace of $\mathbb{F}_{q}^{n}$ of dimension $k$ and the minimum distance between any two vectors in $\mathcal{C}$ is $d$. The generator matrix $G$ of $\mathcal{C}$ is a $k\times n$ matrix such that any set of $k$ columns of $G$ are linearly independent. The standard form of $G$ is $[I|A]$ where $I$ is a $k\times k$ identity matrix and $A$ is $k\times n-k$ matrix. The Singleton bound sets an upper bound on the minimum distance of a code with given dimensions. It says, for an $[n,k,d]$ code, $d\leq n-k+1$. A code that meets this bound is known as a maximum distance separable (MDS) code. An alternative definition of an MDS code, as given in [25], is the following.
@@ 4 Structure of cyclic matrices and their connection with g-circulant matrices
In [1, 13] a detailed study on the properties of $g$-circulant matrices is provided. We first show that we are only interested in the case $\gcd(k,g)=1$. This restriction is vital because if $\gcd(k,g)>1,$ these $g$-circulant matrices never be MDS. To justify this, we use Lemma 3.16.
@@
Let $\gcd(k,g)=d$. Then there exists $k_{1},k_{2}$ such that $k=dk_{1},g=dk_{2}$ and $\gcd(k_{1},k_{2})=1$. In a $g$-circulant matrix, the entries satisfy the relation $a_{i,j}=a_{i+1,j+g}$ for $i,j=0,1,2,\ldots,k-1$, with suffixes calculated modulo $k$. Let us consider the first row of $A$ as $(a_{0,0},a_{0,1},a_{0,2},\ldots,a_{0,n-1})$. According to Lemma 3.16, there exits an integer $\alpha,1<\alpha\leq k-1$ such that $\alpha g=0$. Consequently, we have
@@
Let $A=g$-circulant$(c_{0},c_{1},\ldots,c_{k-1})$ with $\gcd(k,g)=1$. Then $A$ can be expressed as $A=\sum_{i=0}^{k-1}a_{i}Q_{g}P^{i}$, where $P=$ circulant$(0,1,0,\ldots,0)$ and $Q_{g}=g$-circulant$(1,0,0,\ldots,0)$.
@@
Let $\mathfrak{C}_{\rho}(c_{0},c_{1},\ldots,c_{k-1})$ be a cyclic matrix. Then there exists a unique permutation matrix $Q$ such that $\mathfrak{C}Q=$ circulant$(c_{0},c_{\rho(0)},c_{\rho^{2}(0)},c_{\rho^{3}(0)},\ldots,c_{\rho^{k-1}(0)})$.
@@
Let us consider the permutation matrix $Q$ such that it’s $(i,j)$-th entry is given by the rule
@@
Let $\mathfrak{C}Q=\mathcal{C}$. We prove that $\mathcal{C}$ is a circulant matrix by showing that $i$-th row of $\mathcal{C}$ is a right shift of $(i-1)$-th row for $i=0,1,\ldots,k-1$, i.e., $\mathcal{C}(i,j)=\mathcal{C}(i-1,j-1)$ for $11\leq i,j\leq k-$ 1. We will establish this property through induction on $i$.
@@
Calculating the first entry of the $i+1$-th row, we get $\mathcal{C}(i+1,0)=\sum_{l=0}^{k-1}\mathfrak{C}_{\rho}(i+1,l)Q(l,0)=\mathfrak{C}(i+1,0)Q(0,0)=c_{\rho^{-(i+1)}(0)}=c_{\rho^{k-i-1}(0)}=\mathcal{C}(i,k-1)$. This holds because $Q(0,0)=1$ and $Q(l,0)=0$ for $l=2,3,\ldots,k-1$. Similarly, by calculating the other entries of the $i+1$-th row, we get $\mathcal{C}(i+1,j)=\sum_{l=0}^{k-1}\mathcal{C}(i+1,l)Q(l,j)=\mathfrak{C}(i+1,\rho^{j}(0))Q(\rho^{j}(0),j)=c_{\rho^{-(i+1)}(\rho^{j}(0))}=c_{\rho^{-(i+1)+j}(0)}=c_{\rho^{k-(i+1)+j}(0)}=\mathcal{C}(i,j-1)$ for $j=1,2,\ldots,k-1$. Therefore $\mathcal{C}$ is a circulant matrix .
To establish uniqueness, suppose there exists another permutation matrix $Q^{\prime}$ defined using $k$-cycle $\rho^{\prime}$ such that $\mathfrak{C}Q^{\prime}=$ circulant$(c_{0},c_{\rho(0)},c_{\rho^{2}(0)},c_{\rho^{3}(0)},\ldots,c_{\rho^{k-1}(0)})$. This implies $\rho^{i}(0)=\rho^{\prime i}(0)$ for $i=1,\ldots,k-1$. Since both $\rho$ and $\rho^{\prime}$ are $k$-cycle permutation, we can conclude that $\rho=\rho^{\prime}$ and $Q=Q^{\prime}$. ∎
@@
Consider the cyclic matrix $\mathfrak{C}_{\rho}(c_{0},c_{1},c_{2},c_{3},c_{4})$ with $\rho=(0~{}2~{}3~{}1~{}4)$. Then $\mathfrak{C}_{\rho}=\begin{bmatrix}c_{0}&c_{1}&c_{2}&c_{3}&c_{4}\\ c_{4}&c_{3}&c_{0}&c_{2}&c_{1}\\ c_{1}&c_{2}&c_{4}&c_{0}&c_{3}\\ c_{3}&c_{0}&c_{1}&c_{4}&c_{2}\\ c_{2}&c_{4}&c_{3}&c_{1}&c_{1}\\ \end{bmatrix}$. Using the constriction of the permutation matrix given in Theorem 4.3, we have $Q=\begin{bmatrix}1&0&0&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&0&1\\ \end{bmatrix}.$ Therefore $\mathfrak{C}_{\rho}Q$ is the circulant matrix with the first row $(c_{0},c_{2},c_{3},c_{1},c_{4})$. Also, $Q^{-1}=\begin{bmatrix}1&0&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ 0&0&0&0&1\\ \end{bmatrix}=\text{cyclic}_{\rho}(1,0,0,0,0)$.
@@
Let $\mathfrak{C}_{\rho}(c_{0},c_{1},c_{2},\ldots,c_{k-1})$ be a cyclic matrix. Then $\mathfrak{C}_{\rho}=\sum_{i=0}^{k-1}a_{\rho^{i}(0)}P^{i}Q_{\rho}$, where $Q_{\rho}=\text{cyclic}_{\rho}(1,0,0,\ldots,0)$ corresponding to the $k$-cycle $\rho$ and $P=$ circulant$(0,1,0,\ldots,0)$.
@@
Let $\mathfrak{C}_{\rho}=$ cyclic$(c_{0},c_{1},c_{2},\ldots,c_{k-1})$, where $\rho$ is the $k$-cycle permutation $(0~{}g~{}{2g\pmod{k}}~{}{3g\pmod{k}}\cdots{(k-1)g\pmod{k}})$ with $\gcd(g,k)=1$. Then $\mathfrak{C}_{\rho}=\sum_{i=0}^{k-1}a_{i}Q_{g}P^{i}$.
@@ 5 Results on orthogonal g-circulant matrices
To investigate the orthogonal property of $2^{d}\times 2^{d}$ circulant MDS matrices over $\mathbb{F}_{2^{m}}$, Gupta and Ray [20] proved that $A^{2^{d}}$ is a scaler matrix, where $A$ is a circulant matrix. After that, Liu and Sim extended these to left-circulant case in [24]. Since a left-circulant matrix is symmetric, involutory property and orthogonal property of such a matrix are equivalent. Note that orthogonal property is not always preserved under permutation equivalence. Therefore, in this section, we study $g$-circulant matrices, with orthogonal property. In Lemma 5.2, we prove the similar result for $g$-circulant matrices of order $2^{d}\times 2^{d}$ over $\mathbb{F}_{2^{m}}$. For that, we first require the following lemma.
@@
If $g>1$ and $\gcd(g,2^{d})=1$, then $2^{d}$ divides $\frac{g^{2^{d}}-1}{g-1}$.
@@
$\frac{g^{2^{d}}-1}{g-1}=(g+1)(g^{2}+1)(g^{2^{2}}+1)\cdots(g^{2^{d-1}}+1)$. Since $\gcd(g,2^{d})=1$ implies $\gcd(g,2)=1$, i.e., $g$ is a odd number. Therefore $(g+1),(g^{2}+1),\cdots,(g^{2^{d-1}}+1)$ are all even numbers. Hence, each term is divisible by $2$ and there are $d$ such terms. Therefore $2^{d}$ divides the product. ∎
@@
In the case of order $4\times 4$, there are only two distinct $g$-circulant matrices: the circulant and the left-circulant matrices. However, if the order is $22^{d}\times 2^{d},d>$ 2, total $2^{d-1}$ different $g$-circulant matrices exists, corresponding to $g=1,3,5,\ldots,2^{d}-1$. Notably, for $g=1$ we obtain a circulant matrix, and for this, our findings are reduced to the first three results presented in Section $3$ of [20]. On the other hand, for the case $g=2^{d}-1$, resulting matrix is left circulant. In this scenario, Theorem 5.4 provide a more general proof of the results presented in Section $5.2$ of [24].

### v2

# A note on cyclic MDS and non-MDS matrices

@@ 1 Introduction
Symmetric key primitives, like block ciphers and hash functions, rely on various components that provide “confusion and diffusion”. These concepts were introduced by Claude Shannon in $1949$, in his seminal paper “Communication Theory of Secrecy Systems” [32]. Confusion and diffusion layers play a pivotal role in the security and efficiency of cryptographic schemes based on block ciphers and hash functions. The goal of the confusion layer is to hide the statistical relationship between the ciphertext and message whereas, the diffusion layer guarantees that every bit of the message and secret key affects multiple bits of the ciphertext, ensuring that, after several rounds, all output bits depend on every input bit. Therefore, the choice of a diffusion layer plays an important role in providing security against differential cryptanalysis [4] and linear cryptanalysis [25] which are the two most important cryptanalysis of block ciphers. The strength of a diffusion layer is usually assessed by its branch number. As a result, constructing diffusion layers with higher branch numbers and low-cost implementations is a challenge for cipher design. Consequently, many block ciphers use Maximum Distance Separable (MDS) matrices, known for their optimal branch numbers, in their diffusion layers. These matrices are redundant parts of systematic generator matrices of MDS codes and used in popular ciphers like AES [12], LED [17], SQUARE [11], SHARK [34] as well as hash functions such as Whirlpool [3] and PHOTON [16].
The construction of diffusion layer using MDS matrices primarily employs two main approaches: recursive and non-recursive [21]. In recursive construction, a companion or sparse matrix $A$ of order $n\times n$ is used such that $A^{n}$ is an MDS matrix. Block cipher LED [17], and hash function PHOTON [16] used this recursive based construction. In contrast, non-recursive methods encompass various techniques, such as the use of Cauchy matrices [18, 24], Vandermonde matrices [22, 31], Hadamard matrices [33], and their generalizations [28]. Furthermore, inspired by the use of circulant MDS matrix in the diffusion layer of AES, various authors [5, 19, 20, 23] have explored the construction of circulant MDS matrices. In [20], the authors extended circulant matrices to circulant-like and generalized circulant MDS matrices.
@@
In search of orthogonal MDS matrices, Liu and Sim introduced cyclic matrices as a generalization of circulant matrices [23]. This broader class of matrices is expected to exhibit orthogonal or involutory properties, which certain orders of circulant matrices lack over a characteristic $2$ field. While, in [23], it was proved that $2^{d}\times 2^{d}$ left-circulant MDS matrices, a subclass of cyclic matrices, cannot be orthogonal, the orthogonal property for the larger class remained an open question. Furthermore, in [5], the authors introduced $\theta$-circulant matrices over the quasi polynomial ring. Later, in $2022$, Adhiguna et al. studied the orthogonal property of $\theta$-circulant MDS matrices in [2] over the quasi polynomial ring. Several other authors [7, 14] extended the search for MDS matrices over rings and modules. The authors of [29, 30] studied orthogonal MDS matrices over the Galois ring $GR(2^{n},k)$. In $2023,$ Mohsenifar et al. [26] studied the connection between cyclic matrices and generalized Cauchy matrices. They proved that it is possible to construct circulant MDS matrices of orders $3\times 3$ and $5\times 5$ under certain conditions on the entries and the finite fields using generalized Cauchy matrices. Recently, in $2024$, the authors in [27] introduced a direct construction of circulant MDS matrices with easily computable inverses using Vandermonde matrix. They also introduced circulant-like matrices and provided a direct construction method for such matrices.
@@ 3 Notations and Preliminaries
Let $\mathbb{F}_{q}$ denote the finite field with $q=p^{m}$ elements where $p$ is a prime number and $m$ is a positive integer. Consider an $[n,k,d]$ linear error-correcting code $\mathcal{C}$ over $\mathbb{F}_{q}$. This code $\mathcal{C}$ is a subspace of $\mathbb{F}_{q}^{n}$ of dimension $k$ and the minimum distance between any two vectors in $\mathcal{C}$ is $d$. The generator matrix $G$ of $\mathcal{C}$ is a $k\times n$ matrix such that any set of $k$ columns of $G$ are linearly independent. The standard form of $G$ is $[I|A]$ where $I$ is a $k\times k$ identity matrix and $A$ is $k\times(n-k)$ matrix. The Singleton bound sets an upper bound on the minimum distance of a code with given dimensions. It says, for an $[n,k,d]$ code, $d\leq n-k+1$. A code that meets this bound is known as a maximum distance separable (MDS) code. An alternative definition of an MDS code, as given in [24], is the following.
@@ 4 Structure of cyclic matrices and their connection with g-circulant matrices
In [1, 13] a detailed study on the properties of $g$-circulant matrices is provided. We first show that we are only interested in the case $\gcd(k,g)=1$. This restriction is vital because if $\gcd(k,g)>1,$ these $g$-circulant matrices can never be MDS. To justify this, we use Lemma 3.16.
@@
Let $\gcd(k,g)=d$. Then there exists $k_{1},k_{2}$ such that $k=dk_{1},g=dk_{2}$ and $\gcd(k_{1},k_{2})=1$. In a $g$-circulant matrix, the entries satisfy the relation $a_{i,j}=a_{i+1,j+g}$ for $i,j=0,1,2,\ldots,k-1$, with suffixes calculated modulo $k$. Let us consider the first row of $A$ as $(a_{0,0},a_{0,1},a_{0,2},\ldots,a_{0,n-1})$. According to Lemma 3.16, there exists an integer $\alpha,1<\alpha\leq k-1$ such that $\alpha g=0$. Consequently, we have
@@
Let $A=g$-circulant$(c_{0},c_{1},\ldots,c_{k-1})$ with $\gcd(k,g)=1$. Then $A$ can be expressed as $A=\sum_{i=0}^{k-1}c_{i}Q_{g}P^{i}$, where $P=$ circulant$(0,1,0,\ldots,0)$ and $Q_{g}=g$-circulant$(1,0,0,\ldots,0)$.
@@
Let $\mathfrak{C}_{\rho}(c_{0},c_{1},\ldots,c_{k-1})$ be a cyclic matrix. Then there exists a unique permutation matrix $Q$ such that $\mathfrak{C}_{\rho}Q=$ circulant$(c_{0},c_{\rho(0)},c_{\rho^{2}(0)},c_{\rho^{3}(0)},\ldots,c_{\rho^{k-1}(0)})$.
@@
Let us consider the permutation matrix $Q$ such that its $(i,j)$-th entry is given by the rule
@@
Let $\mathfrak{C}_{\rho}Q=\mathcal{C}$. We prove that $\mathcal{C}$ is a circulant matrix by showing that $i$-th row of $\mathcal{C}$ is a right shift of $(i-1)$-th row for $i=0,1,\ldots,k-1$, i.e., $\mathcal{C}(i,j)=\mathcal{C}(i-1,j-1)$ for $11\leq i,j\leq k-$ 1. We will establish this property through induction on $i$.
@@
Calculating the first entry of the $i+1$-th row, we get $\mathcal{C}(i+1,0)=\sum_{l=0}^{k-1}\mathfrak{C}_{\rho}(i+1,l)Q(l,0)=\mathfrak{C}_{\rho}(i+1,0)Q(0,0)=c_{\rho^{-(i+1)}(0)}=c_{\rho^{k-i-1}(0)}=\mathcal{C}(i,k-1)$. This holds because $Q(0,0)=1$ and $Q(l,0)=0$ for $l=2,3,\ldots,k-1$. Similarly, by calculating the other entries of the $i+1$-th row, we get $\mathcal{C}(i+1,j)=\sum_{l=0}^{k-1}\mathfrak{C}_{\rho}(i+1,l)Q(l,j)=\mathfrak{C}_{\rho}(i+1,\rho^{j}(0))Q(\rho^{j}(0),j)=c_{\rho^{-(i+1)}(\rho^{j}(0))}=c_{\rho^{-(i+1)+j}(0)}=c_{\rho^{k-(i+1)+j}(0)}=\mathcal{C}(i,j-1)$ for $j=1,2,\ldots,k-1$. Therefore $\mathcal{C}$ is a circulant matrix .
To establish uniqueness, suppose there exists another permutation matrix $Q^{\prime}$ defined using $k$-cycle $\rho^{\prime}$ such that $\mathfrak{C}_{\rho^{\prime}}Q^{\prime}=$ circulant$(c_{0},c_{\rho(0)},c_{\rho^{2}(0)},c_{\rho^{3}(0)},\ldots,c_{\rho^{k-1}(0)})$. This implies $\rho^{i}(0)=\rho^{\prime i}(0)$ for $i=1,\ldots,k-1$. Since both $\rho$ and $\rho^{\prime}$ are $k$-cycle permutation, we can conclude that $\rho=\rho^{\prime}$ and $Q=Q^{\prime}$. ∎
@@
Consider the cyclic matrix $\mathfrak{C}_{\rho}(c_{0},c_{1},c_{2},c_{3},c_{4})$ with $\rho=(0~{}2~{}3~{}1~{}4)$. Then $\mathfrak{C}_{\rho}=\begin{bmatrix}c_{0}&c_{1}&c_{2}&c_{3}&c_{4}\\ c_{4}&c_{3}&c_{0}&c_{2}&c_{1}\\ c_{1}&c_{2}&c_{4}&c_{0}&c_{3}\\ c_{3}&c_{0}&c_{1}&c_{4}&c_{2}\\ c_{2}&c_{4}&c_{3}&c_{1}&c_{1}\\ \end{bmatrix}$. Using the construction of the permutation matrix given in Theorem 4.3, we have $Q=\begin{bmatrix}1&0&0&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&0&1\\ \end{bmatrix}.$ Therefore $\mathfrak{C}_{\rho}Q$ is the circulant matrix with the first row $(c_{0},c_{2},c_{3},c_{1},c_{4})$. Also, $Q^{-1}=\begin{bmatrix}1&0&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&1&0&0&0\\ 0&0&0&0&1\\ \end{bmatrix}=\text{cyclic}_{\rho}(1,0,0,0,0)$.
@@
Let $\mathfrak{C}_{\rho}(c_{0},c_{1},c_{2},\ldots,c_{k-1})$ be a cyclic matrix. Then $\mathfrak{C}_{\rho}=\sum_{i=0}^{k-1}c_{\rho^{i}(0)}P^{i}Q_{\rho}$, where $Q_{\rho}=\text{cyclic}_{\rho}(1,0,0,\ldots,0)$ corresponding to the $k$-cycle $\rho$ and $P=$ circulant$(0,1,0,\ldots,0)$.
@@
Let $\mathfrak{C}_{\rho}=$ cyclic$(c_{0},c_{1},c_{2},\ldots,c_{k-1})$, where $\rho$ is the $k$-cycle permutation $(0~{}g~{}{2g\pmod{k}}~{}{3g\pmod{k}}\cdots{(k-1)g\pmod{k}})$ with $\gcd(g,k)=1$. Then $\mathfrak{C}_{\rho}=\sum_{i=0}^{k-1}c_{i}Q_{g}P^{i}$.
@@ 5 Results on orthogonal g-circulant matrices
To investigate the orthogonal property of $2^{d}\times 2^{d}$ circulant MDS matrices over $\mathbb{F}_{2^{m}}$, Gupta and Ray [20] proved that $A^{2^{d}}$ is a scalar matrix, where $A$ is a circulant matrix. After that, Liu and Sim extended these to left-circulant case in [23]. Since a left-circulant matrix is symmetric, involutory property and orthogonal property of such a matrix are equivalent. Note that orthogonal property is not always preserved under permutation equivalence. Therefore, in this section, we study $g$-circulant matrices, with orthogonal property. In Lemma 5.2, we prove the similar result for $g$-circulant matrices of order $2^{d}\times 2^{d}$ over $\mathbb{F}_{2^{m}}$. For that, we first require the following lemma.
@@
If $g>1$ be an odd integer then $2^{d}$ divides $\frac{g^{2^{d}}-1}{g-1}$.
@@
$\frac{g^{2^{d}}-1}{g-1}=(g+1)(g^{2}+1)(g^{2^{2}}+1)\cdots(g^{2^{d-1}}+1)$. Since $g$ is an odd number, $(g+1),(g^{2}+1),\cdots,(g^{2^{d-1}}+1)$ are all even numbers. Hence, each term is divisible by $2$ and there are $d$ such terms. Therefore $2^{d}$ divides the product. ∎
@@
In the case of order $4\times 4$, there are only two distinct $g$-circulant matrices: the circulant and the left-circulant matrices. However, if the order is $22^{d}\times 2^{d},d>$ 2, total $2^{d-1}$ different $g$-circulant matrices exists, corresponding to $g=1,3,5,\ldots,2^{d}-1$. Notably, for $g=1$ we obtain a circulant matrix, and for this, our findings are reduced to the first three results presented in Section $3$ of [20]. On the other hand, for the case $g=2^{d}-1$, resulting matrix is left circulant. In this scenario, Theorem 5.4 provides a more general proof of the results presented in Section $5.2$ of [23].